### v1

# Machine Learning and the Physical Sciences

Abstract
Reduced-order models (ROMs) can efficiently simulate high-dimensional physical systems, but lack robust uncertainty quantification methods. Existing approaches are frequently architecture- or training-specific, which limits flexibility and generalization. We introduce a post hoc, model-agnostic framework for predictive uncertainty quantification in latent space ROMs that requires no modification to the underlying architecture or training procedure. Using conformal prediction, our approach estimates statistical prediction intervals for multiple components of the ROM pipeline: latent dynamics, reconstruction, and end-to-end predictions. We demonstrate the method on a latent space dynamical model for cloud microphysics, where it accurately predicts the evolution of droplet-size distributions and quantifies uncertainty across the ROM pipeline.
1 Introduction
Latent space reduced-order modeling learns a compact representation of high-dimensional physical dynamics in a lower-dimensional latent space. These models are valuable for scientific applications where the governing physics are partially known or computationally prohibitive to resolve. For example, accurately resolving clouds and precipitation in an atmospheric simulation would require tracking high-dimensional droplet-size-distributions (DSDs), a longstanding parametric challenge in climate and weather modeling known as “cloud microphysics.” Error and uncertainty in microphysics parameterizations is typically not quantified, but is believed to be a dominant source of uncertainty in future climate projections [morrison_confronting_2020].
Latent space ROMs have proven effective in efficiently simulating related complex fluid mechanical systems (e.g. [choi_sns_2020, fries_lasdi_2022], yet convincing practitioners of their reliability is challenging due to the lack of unified and robust uncertainty quantification (UQ) frameworks. Existing UQ methods for latent space dynamical models are often tied to specific architectures [Bonnet2023, cheng2025latentspaceenergybasedneural, YONG2025117638], require expensive training [Simpson2024-ie], or make parametric assumptions [iakovlev2023latentneuralodessparse, Kingma_2019].
We present a model-agnostic, post hoc framework for predictive UQ in latent-space ROMs that quantifies uncertainty on reconstruction, latent dynamics, and end-to-end predictions without altering the base architecture or training procedure. Our approach utilizes conformal prediction (CP), a distribution-free method that produces statistically valid prediction intervals – a first for latent space ROMs. We demonstrate this UQ pipeline on a cloud microphysics ROM trained to predict the evolution of cloud DSDs during coalescence and the formation of precipitation, producing reliable UQ estimates that allow practitioners to rigorously evaluate individual components of the ROM architecture.
2 Proposed UQ framework
A latent space dynamical ROM consists of a data space, $\mathcal{X}\subseteq\mathbb{R}^{d}$, and a latent space, $\mathcal{Z}\subseteq\mathbb{R}^{D}$, where $D\ll d$; an encoder $E:\mathcal{X}\rightarrow\mathcal{Z}$ and a decoder $D:\mathcal{Z}\rightarrow\mathcal{X}$; and a dynamical system model $F:\mathcal{T}\times\mathcal{Z}\rightarrow\mathcal{Z}$ defined on the latent space, where $\mathcal{T}\subseteq\left[0,T\right]$ for some final time $T>0$. (See Figure 1.)
Figure 1: A generic latent space dynamical model with fixed dynamical time-step $\Delta t$. Figures 1a and 1b show the reconstruction and dynamics sub-models, respectively, that comprise the end-to-end model architecture shown in Figure 1c.
We consider the setting where we observe $n$ independent time-evolving realizations of a physical system in $\mathcal{X}$: $\{x_{t}^{(j)}\}_{t\in\mathcal{T}}$ for $j=1,\dots,n$. The proposed pipeline applies regardless of whether model components are trained separately or end-to-end. Hence, at each fixed time $t\in\mathcal{T}$ and across samples $j=1,\dots,n$, we compute predictive uncertainties in the components of a latent space dynamical model: 1) reconstruction from the latent space, 2) dynamics in the latent space, and 3) the entire pipeline combined (end-to-end), as shown in Figure 1.
Conformal predictions
Conformal prediction (CP) provides statistical prediction intervals by computing nonconformity scores on a held-out calibration dataset such that the true outcome $Y$ is contained in the prediction set $\Gamma(X)$ with probability at least $1-\alpha$ [angelopoulos2021gentle, barber_predictive_2021, steinberger_leave-one-out_2016, vovk_cross-conformal_2015, vovk_algorithmic_2022], i.e., $\mathbb{P}\left(Y\subseteq\Gamma(X)\right)\geq 1-\alpha$. This guarantee holds under the assumption of exchangeability of the calibration and test data, satisfied in the case of our DSD dataset due to independence of the sampled initial conditions—see Appendix C.1 for details. Because CP requires no changes to model architecture or changing or parametric assumptions on the data distribution, it can be applied either to full outputs or component-wise in multi-dimensional settings to obtain reliable, distribution-free guarantees on predictive coverage [angelopoulos2021gentle]. We illustrate three common variants of conformal predictions: Vanilla conformal (train–test split, using training data for scoring discrepancies), split conformal (train–validation–test split, scoring on validation set), and CV+ conformal ($k$-fold train–validation splits with aggregated residuals) [angelopoulos2021gentle, barber_predictive_2021, 9835501]. CV+ generally yields tighter intervals while maintaining coverage guarantees, with the choice of folds $k$ balancing statistical efficiency and computational overhead [barber_predictive_2021].
Rather than assuming errors are symmetric around the predictions, we follow a two-sided approach to construct prediction intervals using the $\alpha/2$ and $1-\alpha/2$ quantiles of the signed residuals [barber_predictive_2021]. This modification produces asymmetric prediction bands with valid $1-\alpha$ coverage—a crucial feature for DSDs, as these are non-negative and often contain values near zero, allowing prediction intervals to reflect physical uncertainty better. See Appendix B for further details.
3 Application to Cloud Microphysics ROM
We demonstrate the UQ pipeline with a cloud microphysics ROM trained on DSDs from particle-based simulations of warm-rain coalescence. Coalescence consists of smaller cloud droplets combining to form larger ones, eventually leading to precipitation [khain_representation_2015]. As shown in Figure 2, droplet mass shifts from smaller to larger size bins under coalescence, and the strong nonlinearity of the process drives the emergence and disappearance of multiple modes.
New parameterizations of cloud microphysics must not only outperform traditional schemes in accuracy and efficiency, but also quantify structural errors and parametric uncertainties that currently hinder the accuracy of large-scale models. Nonlinear ROMs offer a more flexible alternative to bulk parameterizations, which impose restrictive modeling assumptions [kessler_distribution_1969, khairoutdinov_new_2000, seifert_double-moment_2001], and to linear latent space ROMs, which could require inefficiently high-dimensional representations to capture DSD coalescence dynamics accurately [lin_enhanced_2025, romor_non-linear_2023, wan_evolve_2023]. Hence, to model droplet coalescence, we combine an autoencoder (AE) for nonlinear dimensionality reduction with parsimonious latent space ODEs, based on the Sparse Identification of Nonlinear Dynamics (SINDy) technique [brunton_discovering_2016, champion_data-driven_2019]. (For details of the AE-SINDy architecture and training, see Appendix C and the associated code repository. Here, we focus on the UQ pipeline and results.)
Our UQ pipeline is uniquely suited to address both structural uncertainty in the autoencoder compression of DSDs to the latent space and parametric uncertainty in the identified ODEs, tracing how these uncertainties propagate into DSD coalescence predictions. The examples in Figure 2 demonstrate uncertainty for one sample DSD at one time, while Figure 3 aggregates prediction intervals across samples to assess uncertainty arising from different parts of the model as functions of time. Our analysis highlights the ability of the presented UQ pipeline to characterize data-driven ROMs by highlighting the specific scales and processes where model improvements are most necessary.
Figure 2: Initial and final states for three sample DSD trajectories from the dataset, evolved using the AE-SINDy pipeline with empirical prediction intervals provided via CV+ conformal predictions (using $k=20$ folds) at varying nominal coverage levels.
Figure 3: Weighted average prediction interval width vs. time, computed by applying vanilla CP, split CP (using a 6060-2020-2020 train-validation-test split), and CV+ conformal predictions (using $k=20$ folds), respectively, to the indicated components of the AE-SINDy pipeline at varying nominal coverage levels $1-\alpha$. We assess this mean predictive uncertainty by averaging the widths of prediction intervals across samples as a function of time for various miscoverage levels $\alpha$, using integrated totals for DSD predictions and averages for latent variables.
4 Results & Discussion
[…]
Figure 3 illustrates not only predictive accuracy on unseen data but also qualitative patterns of uncertainty propagation across the AE-SINDy architecture. As anticipated, reconstruction uncertainty intervals, which characterize the autoencoder alone, are consistent across time. The latent dynamics exhibit rapid incipient growth in uncertainty that slows over time, reflecting the challenging cloud-to-rain transition before droplets settle into a rain-dominant coalesced state. By contrast, the full ROM produces nearly linear growth in the end-to-end predictive errors. This highlights a key advantage of component-wise uncertainty analysis in ROMs: we detected how latent errors are effectively “smoothed out” by the autoencoder, yielding linear error growth in the final predictions.
Although a dynamical system may evolve on a lower-dimensional manifold, an inaccurate mapping between physical and latent space hinders a faithful, parsimonious latent representation, complicating the modeling of latent dynamics— particularly in the context of SINDy [brunton_discovering_2016, champion_data-driven_2019]. Figure 3 shows that predicted reconstruction errors remain consistent over time for most of the data, regardless of $\alpha$ or the CP method. Even though predicted errors in the latent dynamics increasingly affect the end-to-end model output accuracy as time progresses, the propagation of these dynamics errors is ultimately mediated by reconstruction back to physical space. Thus, for this cloud microphysics ROM, future work to reduce structural uncertainty in the autoencoder will likely have a greater impact on overall model performance than refining the dynamical model.
While all conformal prediction methods achieved near-nominal coverage (see Appendix A), CV+ yielded wider average intervals to produce more reliable coverage, but at a higher computational cost. CV+ requires retraining the surrogate model $k$ times, compared to just once for vanilla or split conformal. The cost of training the surrogate model will therefore determine practical choices for CP techniques in future applications.
5 Limitations & Future Directions
The agreement across CP methods and variability in empirical coverages was notably better for the reconstruction and end-to-end network predictions than for latent dynamical predictions (cf. Figure 3). This is likely due to greater variance in prediction fidelity across times and variables—as well as a wider range of magnitudes overall—in latent predictions compared to normalized DSD predictions. Increasing training data—for instance, by widening the simulation domain used to generate the dataset (see Appendix C.1) or increasing its resolution—could reduce this variability, yielding more consistent CP intervals and improving coverage accuracy on the test set.
While this work demonstrates the flexibility of conformal prediction for uncertainty quantification in general black-box architectures, a key limitation of standard CP methods is that prediction intervals are scaled only relative to the input variables; the width of a given interval at a particular output and time remains fixed across the dataset. Although adaptive variants can adjust interval widths to reflect varying uncertainty [angelopoulos2021gentle, barber2022limits, colombo2023training, guan2023localized, lei2018distribution, romano2019conformalized], we do not explore these extensions in this study.
That being said, the post hoc UQ approach introduced in this work is not limited to conformal prediction and could also extend to other interval- and set-valued UQ methods—e.g., parametric prediction intervals, confidence intervals, and Bayesian credible intervals [casella_statistical_2002, gelman_bayesian_2013, hyndman_forecasting_2021, montgomery_applied_2014, wasserman_all_2004]. Exploring these extensions, especially on other ROMs, could show the usefulness of this approach for quantifying other types of uncertainty in surrogate modeling pipelines.
Acknowledgments and Disclosure of Funding
This work was performed in part under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory (LLNL) under Contract DE-AC52-07NA27344 and supported by the Laboratory Directed Research and Development Program (LDRD), project number 25-ERD-045. J. Katona was partially supported by a stipend and teaching fellowship from the Yale Graduate School of Arts and Sciences. The authors have declared that none of them has any competing interests. Released under IM number LLNL-CONF-2010541.
The authors would like to thank LLNL—particularly Livermore Computing—for their support and for providing high-performance computing resources that enabled the simulations, data analyses, and figure generation in this study.
Appendix A Validation of empirical prediction intervals
Tables 1 and 2 give summary statistics—means and standard deviations, and medians, respectively—for the nominal coverages in the outputs for different subsets of the network: autoencoder/reconstruction alone, latent dynamical model alone, and the entire end-to-end network predictions. These statistics are averaged over all output variables—64 bins for DSD data or 4 latent variables—and times—61 timesteps at 10 second intervals.
Table 1: Empirical coverage (in %, given as empirical mean $\pm$ empirical standard deviation across all times and output coordinates) for prediction intervals at marginal coverage levels 90%, 95%, 98%, and 99%, computed using the CP methods indicated. Split conformal was applied on a 60-20-20 train-validation-test split, and CV+ conformal was applied with $k=20$ folds. Empirical medians can be found in Table 2.
CP Method Sub-model $1-\alpha=90$% $1-\alpha=95$% $1-\alpha=98$% $1-\alpha=99$%
Vanilla Reconstruction 88.56 $\pm$ 3.15 93.85 $\pm$ 2.30 96.98 $\pm$ 1.63 98.01 $\pm$ 1.35
Vanilla Latent dynamics 86.00 $\pm$ 12.70 91.80 $\pm$ 12.66 95.39 $\pm$ 12.57 96.88 $\pm$ 12.64
Vanilla End-to-end 88.62 $\pm$ 3.35 93.71 $\pm$ 2.58 96.79 $\pm$ 1.97 97.93 $\pm$ 1.60
Split Reconstruction 87.70 $\pm$ 4.31 92.87 $\pm$ 3.61 96.10 $\pm$ 2.56 97.34 $\pm$ 2.17
Split Latent dynamics 86.93 $\pm$ 11.80 92.05 $\pm$ 12.22 95.14 $\pm$ 12.44 96.48 $\pm$ 12.54
Split End-to-end 87.38 $\pm$ 4.58 92.78 $\pm$ 3.56 96.26 $\pm$ 2.53 97.47 $\pm$ 1.94
CV+ Reconstruction 89.03 $\pm$ 3.09 94.36 $\pm$ 2.23 97.35 $\pm$ 1.61 98.28 $\pm$ 1.36
CV+ Latent dynamics 91.08 $\pm$ 12.96 95.67 $\pm$ 12.73 97.55 $\pm$ 12.71 98.01 $\pm$ 12.68
CV+ End-to-end 90.47 $\pm$ 3.59 95.22 $\pm$ 2.43 97.66 $\pm$ 1.70 98.49 $\pm$ 1.32
Table 2: Empirical coverage (in %, given as the empirical median across all times and output coordinates) for prediction intervals at marginal coverage levels $1-\alpha=$90%, 95%, 98%, and 99%, computed using the CP methods indicated. Split conformal was applied on a 60-20-20 train-validation-test split, and CV+ conformal was applied with $k=20$ folds. Empirical means and empirical standard deviations can be found in Table 1.
CP Method Sub-model $1-\alpha=90$% $1-\alpha=95$% $1-\alpha=98$% $1-\alpha=99$%
Vanilla Reconstruction 88.71 93.55 96.77 98.39
Vanilla Latent dynamics 89.52 94.35 96.77 99.19
Vanilla End-to-end 88.71 93.55 97.58 98.39
Split Reconstruction 87.10 93.55 95.97 97.58
Split Latent dynamics 88.71 94.35 96.77 98.39
Split End-to-end 87.10 92.74 96.77 97.58
CV+ Reconstruction 89.11 94.35 97.58 98.39
CV+ Latent dynamics 93.55 97.58 100.00 100.00
CV+ End-to-end 91.13 95.16 97.58 99.19
Across all CP methods, the empirical coverage was generally close to nominal levels, indicating good calibration. While split conformal is theoretically more accurate than vanilla conformal [angelopoulos2021gentle, lei2018distribution], for both reconstruction and end-to-end outputs, vanilla and split conformal performed similarly, with mean coverages typically within 1%–2% of the target and relatively low variability and standard deviations under $\sim$4%. By contrast, CV+ consistently achieved slightly higher accuracy, producing empirical coverages that were closer to the nominal rates in both mean and median, especially at higher confidence levels. This improvement was most apparent in the end-to-end model, where CV+ coverage levels tracked the nominal ones more tightly than vanilla or split conformal did.
The latent dynamical model posed the greatest challenge for all CP methods, with notably higher variability in empirical coverage. Standard deviations exceeded 12% for vanilla and split conformal across all levels, reflecting unstable coverage across output variables and times. That being said, CV+ was partially able to mitigate this issue, yielding higher mean and median coverages that sometimes matched or exceeded the target levels. For instance, CV+ displayed a median coverage of 93.55% vs. 90% at the 90% nominal level, and even full coverage at the 98% and 99% levels. These results suggest that, while CP methods reliably capture uncertainty for reconstructions and end-to-end outputs, consistently quantifying predictive uncertainty for latent dynamics remains challenging, though CV+ provides the most robust coverage.
To test more specifically for consistency in the predictive intervals across different CP methods, we also refer the reader to Figure 3. For all three subsets of the network, the predictive errors became less consistent as $\alpha\rightarrow 0$. This is because the convergence of empirical quantiles to their true values depends strongly on the miscoverage rate $\alpha$ [serfling1980chapter2, vaart1998chapter21]. In particular, for CP, predictive intervals at smaller $\alpha$ require larger calibration sets to stabilize because extreme quantiles converge more slowly, whereas more central quantiles yield more stable intervals with fewer samples [lei2018distribution, vovk_algorithmic_2022].
Appendix B Tailwise conformal prediction intervals
In the more basic forms of conformal prediction, uncertainty intervals are constructed symmetrically—predictive errors above and below are assumed to have the same distribution. Concretely, for a model $f:X\rightarrow Y$, if we denote the residual
R:=yf(x)R:=y-f(x)
between the model prediction $f(x)$ for an input $x\in X$ and a true outcome $y\in Y$, then the absolute residuals $|R|$ are often used to calibrate a single quantile. Thus, the prediction interval is symmetric and can be written as
$\left[f(x)-Q_{1-\alpha}(|R|),f(x)+Q_{1-\alpha}\left(|R|\right)\right],$
where $Q_{1-\alpha}\left(|R|\right)$ is the $(1-\alpha)$-quantile of $|R|$ as computed empirically over the dataset.
The aforementioned construction is simple and guarantees the desired coverage over the calibration data, but it forces the lower and upper bounds to be equally wide. Hence, following the tailwise quantile approach first introduced in [barber_predictive_2021], we use tailwise (one-sided) quantiles. For a desired miscoverage $\alpha$, we split the miscoverage evenly between the tails on both sides, with $\alpha/2$ mass for each. We then compute the lower and upper quantiles for the signed residuals: $Q_{\alpha/2}(R)$ and $Q_{1-\alpha/2}(R)$. This defines the prediction intervals used in this study:
$\left[f(x)+Q_{\alpha/2}(R),f(x)+Q_{1-\alpha/2}(R)\right],$
which still guarantees $1-\alpha$ coverage over the validation set but allows the lower and upper margins to differ whenever the residual distribution is asymmetric.
Appendix C Reproducibility Details for AE-SINDy Model Training
This appendix provides the essential information required to reproduce the autoencoder-SINDy (AE-SINDy) model setup and training procedure used in this study. The description covers data acquisition and preprocessing, model architecture and hyperparameters, and the training workflow, including loss function specification. The source code and data are further included in a linked repository.
C.1 Data Source and Preprocessing
The AE-SINDy model is trained using binned particle size distribution (PSD) data generated from a large-eddy simulation (LES) employing the superdroplet method. We simulate the evolution of a warm liquid-phase cloud that forms from a gaussian surface moisture and heat flux, growing in altitude before precipitating. The primary datasets are accessed in NetCDF format and contain variables for binned droplet mass distributions over time and space with droplet coalescence active as the only enable droplet dynamic. For model input, only samples with sufficient liquid water content (e.g., exceeding $10^{-5}$ kg/kg) are included. Each PSD is normalized by its total liquid mass to ensure scale invariance during encoding.
The dataset is partitioned into training (80%, 494 samples) and testing (20%, 124 samples) sets. For each sample, the normalized PSD and its time derivative (computed via finite differences) are paired with the corresponding total liquid mass. The input tensors are shaped as $(N_{\text{batch}},N_{\text{bins}})$ for both the PSD and its time derivative, where $N_{\text{bins}}=64$. The total mass is provided as an additional input feature that bypasses the encoder to become the final latent variable. Total mass is further rescaled during training and testing by the maximum value of total mass contained in the training dataset.
C.2 Model Architecture and Hyperparameters
The AE-SINDy model consists of three main components:
  1. 1.
    Encoder: A feed-forward neural network (FFNN) with four fully-connected layers, sequentially reducing the input dimension from $N_{\text{bins}}$ to the latent dimension (minus the total mass variable), i.e. $N_{\text{latent}}-1=3$. ReLU activations are used on hidden layers, with the final layer mapping to the latent space without nonlinearity.
  2. 2.
    Decoder: A FFNN mirroring the encoder structure, with four fully-connected layers expanding from $N_{\text{latent}}-1$ back to $N_{\text{bins}}$. The output layer applies a softmax activation to ensure the reconstructed PSD is normalized.
  3. 3.
    SINDy Dynamics Module: A bias-free, single-layer neural network representing the sparse identification of nonlinear dynamics (SINDy). This layer outputs a linear combination of time derivatives in the latent space using a polynomial feature library up to order 2.
Key hyperparameters for the model are as follows:
  • latent_dim: Number of latent variables (4 total: 3 from the PSD encoding, plus 1 for the total liquid mass)
  • poly_order: Maximum polynomial order in SINDy library (2)
  • batch_size: Training batch size (25)
  • learning_rate: Initial learning rate for AdamW optimizer (e.g., 0.0042)
  • patience: Early stopping patience (50 epochs)
  • weight_decay: L2 regularization coefficient ($10^{-3}$)
  • tol: Numerical tolerance for loss calculations ($10^{-8}$)
  • loss_weights: Relative weights for loss terms, determined via Champion et al.’s recommended scaling (see code for details)
All network weights are initialized using Xavier or Kaiming normal initialization, with zero bias.
C.3 Training Procedure and Loss Function
Training is performed using the AdamW optimizer with learning rate scheduling and early stopping based on validation loss. The model is trained for up to 1000 epochs, with the option to halt training if no improvement is observed over a specified patience interval.
The total loss function $L$ is a weighted sum of three components:
$L=L_{\text{recon}}+w_{dx}L_{dx}+w_{dz}L_{dz}$ (1)
where:
  • $L_{\text{recon}}$ is the Kullback-Leibler divergence between the normalized input PSD and its reconstruction.
  • $L_{dx}$ is the mean squared error between the predicted and actual time derivative of the PSD, projected via the decoder.
  • $L_{dz}$ is the mean squared error between the predicted and actual time derivative in the latent space, as computed by the SINDy module.
Loss weights are chosen to balance the reconstruction and dynamics learning, following a scaling based on the relative magnitudes of the PSD and its time derivative in the training data as in [champion_data-driven_2019]. Other parameters, including the batch size, initial learning rate, and a multiplicative factor of $w_{dx}$, were determined using hyperparameter optimization.
C.4 Code Availability
All code used for data processing, model definition, and training is written in Python using PyTorch and is available at https://anonymous.4open.science/r/UQ_AE-SINDy-1A36. The scripts include utilities for loading NetCDF datasets, constructing PyTorch DataLoaders, defining the AE-SINDy architecture, executing the training loop, and running and visualizing the uncertainty quantification pipelines.